Write a program to explain the how the intensity varies as a function of phase difference / path difference in interference of two sinusoidal waves.
1
Expert's answer
2019-10-16T09:20:44-0400

The intensity of the net waves disturbance can be expressed

"I=4I_0 \cos^2 {\frac{δ}{2}} × \cos^2 {(ѡt+\frac{δ}{2})}"


where δ is the phase difference


We can rewrite this equation as

"I=2I_0 \cos^2 {\frac{δ}{2}}"